# Learning Geodesics of Geometric Shape Deformations From Images

### v1

@@ Abstract
This paper presents a novel method, named geodesic deformable networks (GDN), that for the first time enables the learning of geodesic flows of deformation fields derived from images. In particular, the capability of our proposed GDN being able to predict geodesics is important for quantifying and comparing deformable shape presented in images. The geodesic deformations, also known as optimal transformations that align pairwise images, are often parameterized by a time sequence of smooth vector fields governed by nonlinear differential equations. A bountiful literature has been focusing on learning the initial conditions (e.g., initial velocity fields) based on registration networks. However, the definition of geodesics central to deformation-based shape analysis is blind to the networks. To address this problem, we carefully develop an efficient neural operator to treat the geodesics as unknown mapping functions learned from the latent deformation spaces. A composition of integral operators and smooth activation functions is then formulated to effectively approximate such mappings. In contrast to previous works, our GDN jointly optimizes a newly defined geodesic loss, which adds additional benefits to promote the network regularizability and generalizability. We demonstrate the effectiveness of GDN on both 2D synthetic data and 3D real brain magnetic resonance imaging (MRI).
Keywords: Geodesic deformations, Diffeomorphisms, Neural Operator
@@ 1 Introduction
Deformable shape provides important features of objects presented in images, for example, abnormal shape changes of anatomical brain structures are shown to be predictors of neurodegenerative disorders (Qiu et al., 2008; Lorenzi et al., 2010; Wang et al., 2007). Existing methods have studied a variety of shape representations, including deformation-based descriptors that focus on highly detailed geometric information on dense images (Christensen et al., 1993; Rueckert et al., 2003). With the underlying assumption that objects in many generic classes can be described as deformed versions of an ideal template, descriptors in this class arise naturally by aligning the template to an input image. The resulting transformation between images is then considered as a shape descriptor that reflects local geometric changes. A rich body of literature has been dedicated to harnessing such shape information for enhanced performance in image analysis, particularly in specialized domains of medical image analysis (Bao et al., 2019; Beekman et al., 2022; Sun et al., 2022) and computational anatomy (Wang et al., 2007; Qiu et al., 2008; Grenander and Miller, 1998). The realm of related research involves a suite of tools designed for comparing, matching, and modeling shapes based on distance metrics in deformation spaces (Qiu et al., 2008; Hong et al., 2017; Niethammer et al., 2019; Qiu et al., 2012). A geodesic (a.k.a., the shortest distance) between images that quantifies object geometric similarity serves as a solid mathematical foundation for deformable shape analysis, including regression (Hong et al., 2017; Niethammer et al., 2011), longitudinal analysis (Singh et al., 2013; Hong et al., 2019), and group comparisons (Miller, 2004; Qiu et al., 2008). Such a metric was first developed in the large deformation diffeomorphic metric mapping (LDDMM) algorithm (Beg et al., 2005; Vialard et al., 2012a), giving rise to a variational principle that expresses the optimal deformation as a geodesic flow of diffeomorphisms (i.e., bijective, smooth, and invertible smooth mappings). The geodesic path of deforming one image to another also serves to regularize the smoothness of the transformations grids with an intact topology of objects.
@@
Recent advances in deep registration networks have achieved remarkable success in learning diffeomorphic transformations between images (Yang et al., 2017; Balakrishnan et al., 2019; Wang and Zhang, 2020; Chen et al., 2022; Kim et al., 2022). These learning-based approaches characterize deformation fields using a time sequence of smooth vector fields governed by nonlinear differential equations. In contrast to traditional registration methods such as LDDMM that rely on iterative optimization schemes (Beg et al., 2005; Vialard et al., 2012a), learning-based methods offer rapid predictions without the need for adhoc manual parameter tuning during testing phases. Despite the aforementioned advantages, current registration networks primarily focus on learning the initial conditions of deformation models, i.e., initial velocity fields (Yang et al., 2017; Wang and Zhang, 2020; Hinkle et al., 2018). The final deformation field is then obtained by solving a set of differential equations, which can be computationally expensive in high-dimensional image space during testing. More importantly, the definition of geodesic metrics in the deformation spaces, which is critical for quantifying and analyzing shape changes, is not explicitly considered or overlooked in the network learning process. This limits the interpretability and regularizability of the transformation process. Additionally, it may also negatively impact the model generalizability by elevating the risk of overfitting the registration loss during training inference.
@@
Recent research on neural operators have been explored to learn the solution of orginary /partical differential equations (O/PDEs) (Lu et al., 2019; Bhattacharya et al., 2021; Li et al., 2020b, a). Among these, a notable work is the Fourier neural operator (FNO) (Li et al., 2020a), which achieves an inference time orders of magnitude faster than traditional numerical solutions to the fluid flow Navier-Stokes equations (Constantin, 1988), without sacrificing accuracy. Inspired by the work of (Li et al., 2020a), a research group had an initial attempt to utilize Fourier neural operators to as a surrogate to learn the solution to geodesic flows governed by the well known Euler-Poincaré differential equation (EPDiff) (Arnold, 1966; Miller et al., 2006) in the context of optimization-based image registration (Wu and Zhang, 2023). However, it still needs a conventional iterative optimization-based scheme for estimating the initial velocity fields during the registration inference phase.
@@
In this paper, we introduce a novel geodesic deformable networks (GDN) that for the first time learn the geodesic metrics of latent deformation models directly from training images in the context of registration network, and ground our backbone of geodesic learning on FNO in the latent space. In contrast to the recent work NeurEPDiff (Wu and Zhang, 2023), our model differs by treating the geodesics as unknown mapping functions in the latent deformation spaces, which are explicitly learned in the context of deep registration networks, without requiring optimization during the registration reference stage. A composition of integral operators and smooth activation functions is then formulated to effectively approximate such mappings in each hidden layer of the proposed GDNs. Our network will jointly optimize a newly defined geodesic loss, guided by the underlying transformation process solved through numerical integrators (Vialard et al., 2012a; Beg et al., 2005). A major benefit of the learned geodesic mapping function lies in its application as a computationally efficient surrogate model for the original numerical solutions of differential equations required to generate final deformations. In contrast to previous works (Balakrishnan et al., 2019; Wang and Zhang, 2020; Chen et al., 2022; Kim et al., 2022), the advantages of our proposed method, GDN, are threefold:
  1. (i)
@@
    Predict a geodesic flow of deformation fields that are important for deep nets to quantify and analyze deformation-based shape objects.
  2. (ii)
@@
  3. (iii)
    Eliminate the need of numerical solutions of differential equations for final deformations in testing inference, a strategy that can be computationally advantageous in high-dimensional image spaces.
Experiments on 2D synthetic data and 3D real brain MRIs show that GDN is able to predict fairly close geodesics compared to conventional numerical solutions. We also demonstrate that GDN achieves superior performance in regularizability and generalizability when comparing to the state-of-the-art deep learning-based registration networks in out-of-distribution (OOD) data (Balakrishnan et al., 2019; Hinkle et al., 2018; Chen et al., 2022; Yang et al., 2017).
@@ 2.1 Geodesics of Diffeomorphisms
Let $\mathrm{Diff}^{\infty}(\Omega)$ denote the space of smooth diffeomorphisms on an image domain $\Omega$. The tangent space of diffeomorphisms is the space $V=\mathfrak{X}^{\infty}(T\Omega)$ of smooth vector fields on $\Omega$. Consider a time-varying velocity field, $\{v_{t}\}:[0,\tau]\rightarrow V$, we can generate diffeomorphisms $\{\phi_{t}\}$ as a solution to the equation
@@
[…]
According to a well known geodesic shooting algorithm (Vialard et al., 2012a), the minimum of the previously defined distance metric above is uniquely determined by solving a Euler-Poincaré differential (EPDiff) equation (Vialard et al., 2012a; Younes et al., 2009) with a given initial condition. That is, for $\forall v_{0}\in V$, at $t=0$, a geodesic path $t\mapsto\phi_{t}\in\mathrm{Diff}^{\infty}(\Omega)$ in the space of diffeomorphisms can be computed by forward shooting the EPDiff equation formulated as
@@ 2.2 LDDMM To Derive Geodesics From Images
Assume a source image $S$ and a target image $T$ defined on a $d$-dimensional torus domain $\Omega=\mathbb{R}^{d}/\mathbb{Z}^{d}$ ($S(x),T(x):\Omega\rightarrow\mathbb{R}$). The space of diffeomorphisms is denoted by $\mathrm{Diff}(\Omega)$. The problem of diffeomorphic image registration is to find the shortest path, i.e., geodesic, to generate time-varying diffeomorphisms $\{\phi_{t}\}:t\in[0,\tau]$, such that a deformed source image by the smooth mapping $\phi_{\tau}$, noted as $S(\phi_{\tau})$, is similar to $T$. By parameterizing the transformation $\phi_{\tau}$ with an initial velocity field $v_{0}$, we write the optimization of diffeomorphic registration in the setting of LDDMM with geodesic shooting as
@@
Here, Dist(·,·) is a distance function that measures the dissimilarity between images and $\lambda$ is a positive weighting parameter. The commonly used distance functions include the sum-of-squared intensity differences ($L_{2}$-norm) (Beg et al., 2005), normalized cross correlation (NCC) (Avants et al., 2008), and mutual information (MI) (Wells et al., 1996). In this paper, we will use the sum-of-squared intensity differences.
@@ 3.1 Geodesic Neural Operator
  1. (i)
    An encoder $P$: lift the input latent feature, $z_{v_{0}}$, to a higher dimensional representation with an increased number of channels.
  2. (ii)
    A geodesic evolution layer: learn a nonlinear mapping function of $z_{v_{t}}\mapsto z_{v_{t+1}}$ along the geodesic path.
  3. (iii)
    A decoder $Q$: project the latent mapping function back to the original input dimension aligned with $z_{v_{0}}$.
Analogous to (Li et al., 2020a), the encoder $P$ and decoder $Q$ can be local transformations parameterized by a shallow fully connected neural network. Please note that the functionalities of $P$ and $Q$ slightly differ from a typical image encoder and decoder, as they expand the latent features, $z_{0}$, into higher-dimensional channels and then project them back to the original dimension of channels. This particular design aims to enrich the model capacity and expressivity in the latent space. Next, we will introduce details of the new geodesic evolution layer.
@@
We utilize a multilayer neural network with $J$ hidden layers to simulate the geodesic mapping function from $z_{v_{t}}$ to $z_{v_{t+1}}$. To simplify the math notation, we let $u^{j}$ denote the input of $j$-th hidden layer and $u^{j+1}$ denotes the output, where $j\in\{1,\cdots,J\}$. A composition of a local linear transformation, $W^{j}$, and a global convolutional kernel, $\mathcal{H}^{j}$, is used to extract both global and local representations at each hidden layer. Additionally, a nonlinear activation function, $\sigma(\cdot)$, is carefully designed to encourage a smooth evolution of the geodesic path. We update $u^{j+1}\mapsto u^{j}$ by
@@
where * represents a convolution operator. In order to compute the global convolution in an efficient way, we apply matrix-vector multiplication for two fourier signals in the frequency domain, and then project it back to the original space. We employ a Gaussian error linear unit with the smoothing operator $K$ to ensure the smoothness of the output signal, we have $\sigma:=K(\text{GeLU}(\cdot))$.
Finally, we update the $z_{v_{t-1}}$ to $z_{v_{t}}$ by
@@
We formulate the geodesic loss function of GNO as an empirical data loss guided by numerical solutions (e.g., using an Euler integrator) to the geodesic shooting equations. Since our model GNO is now learning a sequence of latent velocity fields $\{z_{v_{0}},z_{v_{1}},\cdots,z_{v_{\tau}}\}$, the registration network will decode a time-sequence of $\{v_{0},v_{1},\cdots,v_{\tau}\}$ accordingly.
Given a set of $N$ observations, $\bm{\hat{v}}^{n}\triangleq\{\hat{v}_{1}^{n},\cdots,\hat{v}_{\tau}^{n}\}$, which are numerical solutions to the geodesic shooting equation (Eq. (2)) with a given initial condition. Our geodesic loss can be formulated as
$l(\theta_{r},\theta_{v})=\sum_{n=1}^{N}\|\mathcal{D}_{\theta_{v}}(\mathcal{G}_{\theta_{r}}(z_{v_{0}}))-\bm{\hat{v}}^{n}\|_{2}^{2}+\text{Reg}(\theta_{r},\theta_{v}),$ (4)
where $\mathcal{D}_{\theta_{v}}(\cdot)$ denotes a registration network decoder, and $\text{Reg}(\cdot)$ is a regularization on the network parameters.

### v3

@@ Abstract
This paper presents a novel method, named geodesic deformable networks (GDN), that for the first time enables the learning of geodesic flows of deformation fields derived from images. In particular, the capability of our proposed GDN being able to predict geodesics is important for quantifying and comparing deformable shape presented in images. The geodesic deformations, also known as optimal transformations that align pairwise images, are often parameterized by a time sequence of smooth vector fields governed by nonlinear differential equations. A bountiful literature has been focusing on learning the initial conditions (e.g., initial velocity fields) based on registration networks. However, the definition of geodesics central to deformation-based shape analysis is blind to the networks. To address this problem, we carefully develop an efficient neural operator to treat the geodesics as unknown mapping functions learned from the latent deformation spaces. A composition of integral operators and smooth activation functions is then formulated to effectively approximate such mappings. In contrast to previous works, our GDN jointly optimizes a newly defined geodesic loss, which adds additional benefits to promote the network regularizability and generalizability. We demonstrate the effectiveness of GDN on both 2D synthetic data and 3D real brain magnetic resonance imaging (MRI). Our code is publicly available at https://github.com/nellie689/GDN.
keywords:
Geodesic deformations, Diffeomorphisms, Neural Operator
doi:
10.59275/j.melba.2025-f9f4
volume: 3
@@ 1 Introduction
Deformable shape provides important features of objects presented in images, for example, abnormal shape changes of anatomical brain structures are shown to be predictors of neurodegenerative disorders (Qiu et al., 2008; Lorenzi et al., 2010; Wang et al., 2007; Wang and Zhang, 2022). Existing methods have studied a variety of shape representations, including deformation-based descriptors that focus on highly detailed geometric information on dense images (Christensen et al., 1993; Rueckert et al., 2003). With the underlying assumption that objects in many generic classes can be described as deformed versions of an ideal template, descriptors in this class arise naturally by aligning the template to an input image. The resulting transformation between images is then considered as a shape descriptor that reflects local geometric changes. A rich body of literature has been dedicated to harnessing such shape information for enhanced performance in image analysis, particularly in specialized domains of medical image analysis (Bao et al., 2019; Beekman et al., 2022; Sun et al., 2022) and computational anatomy (Wang et al., 2007; Qiu et al., 2008; Grenander and Miller, 1998). The realm of related research involves a suite of tools designed for comparing, matching, and modeling shapes based on distance metrics in deformation spaces (Qiu et al., 2008; Hong et al., 2017; Niethammer et al., 2019; Qiu et al., 2012). A geodesic (a.k.a. a curve whose tangent vector is parallel transported along itself) locally minimizes the Riemannian distance between points on a manifold. Such geodesics can quantify the geometric similarity between objects and serve as a solid mathematical foundation for deformable shape analysis, including regression (Hong et al., 2017; Niethammer et al., 2011), longitudinal analysis (Singh et al., 2013; Hong et al., 2019), and group comparisons (Miller, 2004; Qiu et al., 2008). Such a metric was first developed in the large deformation diffeomorphic metric mapping (LDDMM) algorithm (Beg et al., 2005; Vialard et al., 2012a), giving rise to a variational principle that expresses the optimal deformation as a geodesic flow of diffeomorphisms (i.e., bijective, smooth, and invertible smooth mappings). The geodesic path of deforming one image to another also serves to regularize the smoothness of the transformations grids with an intact topology of objects.
@@
Recent advances in deep registration networks have achieved remarkable success in learning diffeomorphic transformations between images (Yang et al., 2017; Balakrishnan et al., 2019; Wang and Zhang, 2020; Chen et al., 2022; Kim et al., 2022). These learning-based approaches characterize deformation fields using a time sequence of smooth vector fields governed by nonlinear differential equations. In contrast to traditional registration methods such as LDDMM that rely on iterative optimization schemes (Beg et al., 2005; Vialard et al., 2012a), learning-based methods offer rapid predictions without the need for adhoc manual parameter tuning during the testing phase. Despite the aforementioned advantages, current registration networks primarily focus on learning the initial conditions of deformation models, i.e., initial velocity fields (Yang et al., 2017; Wang and Zhang, 2020; Hinkle et al., 2018). The final deformation field is then obtained by solving a set of differential equations, which can be computationally expensive in high-dimensional image space during testing. More importantly, the definition of geodesic metrics in the deformation spaces, which is critical for quantifying and analyzing shape changes, is not explicitly considered or overlooked in the network learning process. This limits the interpretability and regularizability of the transformation process. Additionally, it may negatively impact the model generalizability by elevating the risk of overfitting the registration loss during training phase.
@@
Recent research on neural operators have been explored to learn the solution of ordinary /partial differential equations (O/PDEs) (Lu et al., 2019; Bhattacharya et al., 2021; Li et al., 2020b, a). Among these, a notable work is the Fourier neural operator (FNO) (Li et al., 2020a), which achieves an inference time orders of magnitude faster than traditional numerical solutions to the fluid flow Navier-Stokes equations (Constantin, 1988), without sacrificing accuracy. Inspired by the work of (Li et al., 2020a), a research group had an initial attempt to utilize Fourier neural operators as a surrogate to learn the solution to geodesic flows governed by the well-known Euler-Poincaré differential equation (EPDiff) (Arnold, 1966; Miller et al., 2006) in the context of optimization-based image registration (Wu and Zhang, 2023). Similarly, a group of research work formulate image registration as learning a transformation flow between images, where the deformation trajectories are obtained by integrating time-dependent velocity fields approximated by neural networks within a neural ODE framework (Sun et al., 2024; Wu et al., 2022). However, they still need a conventional iterative optimization-based scheme for estimating the initial velocity fields during the registration inference phase.
@@
In this paper, we introduce a novel geodesic deformable network (GDN) that for the first time learns the geodesic metrics of latent deformation models directly from training images in the context of registration networks. Our method builds upon FNO Li et al. (2023) to model geodesic learning in a latent space, enabling efficient approximation of complex deformation dynamics. Unlike existing approaches (Sun et al., 2024; Wu et al., 2022), our proposed method takes a fundamentally different approach by leveraging the EPDiff geodesic shooting equation to guide the learning of time-dependent velocity fields constrained along geodesic paths. In comparison to the recent NeurEPDiff method (Wu and Zhang, 2023), our model differs by treating the geodesics as unknown mapping functions in the latent deformation spaces, which are explicitly learned in the context of deep registration networks, without requiring optimization during the registration inference stage. A composition of integral operators and smooth activation functions is then formulated to effectively approximate such mappings in each hidden layer of the proposed GDNs. Our network will jointly optimize a newly defined geodesic loss, guided by the underlying transformation process solved through numerical integrators (Vialard et al., 2012a; Beg et al., 2005). A major benefit of the learned geodesic mapping function lies in its application as a computationally efficient surrogate model for the original numerical solutions of differential equations required to generate final deformations. In contrast to previous work (Balakrishnan et al., 2019; Wang and Zhang, 2020; Chen et al., 2022; Kim et al., 2022), the advantages of our proposed method, GDN, are threefold:
@@
  • Predict a geodesic flow of deformation fields using a surrogate model, which is important for deep nets to quantify and analyze deformation-based shape objects.
@@
  • Eliminate the need for numerical solutions of differential equations for final deformations in testing inference, a strategy that can be computationally advantageous in high-dimensional image spaces.
Experiments on 2D synthetic data and 3D real brain MRIs show that GDN is able to predict fairly close geodesics compared to conventional numerical solutions. We also demonstrate that GDN achieves superior performance in regularizability and generalizability when compared to the state-of-the-art deep learning-based registration networks in out-of-distribution (OOD) data (Balakrishnan et al., 2019; Hinkle et al., 2018; Chen et al., 2022; Yang et al., 2017).
@@ 2.1 Geodesics of Diffeomorphisms
Let $\mathrm{Diff}^{\infty}(\Omega)$ denote the space of smooth diffeomorphisms on an image domain $\Omega$. The tangent space of diffeomorphisms is the space $V=\mathfrak{X}^{\infty}(T\Omega)$ of smooth vector fields on $\Omega$. Consider a time-varying velocity field, $\{v_{t}\}:[0,\tau]\rightarrow V$, we can then generate diffeomorphisms $\{\phi_{t}\}$ as a solution to the equation
@@
The LDDMM (Beg et al., 2005) provides a distance metric in the space of diffeomorphisms, which is also used as a regularization for image registration. Such a distance metric is formulated as an integral of the Sobolev norm of the time-dependent velocity field $v_{t}$, i.e., $\int_{0}^{\tau}(\mathcal{L}v_{t},v_{t})\,dt$, where $\mathcal{L}:V\rightarrow V^{*}$ is a symmetric, positive-definite differential operator that maps a tangent vector $v(t)\in V$ into its dual space as a momentum vector $m(t)\in V^{*}$. We typically write $m(t)=\mathcal{L}v(t)$, or $v(t)=\mathcal{K}m(t)$, with $\mathcal{K}$ being an inverse operator of $\mathcal{L}$. In this paper, we adopt a commonly used Laplacian operator $\mathcal{L}=(-\alpha\Delta+\beta\,\text{Id})^{2}$, where $\alpha$ and $\beta$ are weighting parameters that control the smoothness of transformation fields and Id is an identity matrix. The $(\cdot,\cdot)$ is a dual paring, which is similar to an inner product between vectors.
Figure 1: An overview of our proposed network GDN.
According to a well-known geodesic shooting algorithm (Vialard et al., 2012a), the minimum of the previously defined distance metric above is uniquely determined by solving an Euler-Poincaré differential (EPDiff) equation (Vialard et al., 2012a; Younes et al., 2009) with a given initial condition. That is, for $\forall v_{0}\in V$, at $t=0$, a geodesic path $t\mapsto\phi_{t}\in\mathrm{Diff}^{\infty}(\Omega)$ in the space of diffeomorphisms can be computed by forward shooting the EPDiff equation formulated as
@@ 2.2 LDDMM To Derive Geodesics From Images
Assume a source image $S$ and a target image $T$ defined on a $d$-dimensional torus domain $\Omega=\mathbb{R}^{d}/\mathbb{Z}^{d}$ ($S(x),T(x):\Omega\rightarrow\mathbb{R}$). The space of diffeomorphisms is denoted by $\mathrm{Diff}(\Omega)$. The problem of diffeomorphic image registration is to find the geodesic, to generate time-varying diffeomorphisms $\{\phi_{t}\}:t\in[0,\tau]$, such that a deformed source image by the smooth mapping $\phi_{\tau}$, noted as $S(\phi_{\tau})$, is similar to $T$. By parameterizing the transformation $\phi_{\tau}$ with an initial velocity field $v_{0}$, we write the optimization problem of diffeomorphic registration in the setting of LDDMM with geodesic shooting as
$E(v_{0})=(\mathcal{L}v_{0},v_{0})+\lambda\text{Dist}(S(\phi_{\tau}),T)$
@@
Here, Dist(·,·) is a distance function that measures the dissimilarity between images and $\lambda$ is a positive weighting parameter. The commonly used distance functions include the sum-of-squared intensity differences ($L_{2}$-norm) (Beg et al., 2005), normalized cross correlation (NCC) (Avants et al., 2008), and mutual information (MI) (Wells et al., 1996). In this paper, we use the sum-of-squared intensity differences.
@@ 3.1 Geodesic Neural Operator
  • An encoder $P$: lifts the input latent feature, $z_{v_{0}}$, to a higher dimensional representation with an increased number of channels.
  • A geodesic evolution layer: learns a nonlinear mapping function of $z_{v_{t}}\mapsto z_{v_{t+1}}$ along the geodesic path.
  • A decoder $Q$: projects the latent mapping function back to the original input dimension aligned with $z_{v_{0}}$.
Analogous to (Li et al., 2020a), the encoder $P$ and decoder $Q$ can be local transformations parameterized by a shallow fully connected neural network. Please note that the functionalities of $P$ and $Q$ slightly differ from a typical image encoder and decoder, as they expand the latent features, $z_{0}$, into higher-dimensional channels and then project them back to the original channel dimension. This particular design aims to enrich the model capacity and expressivity in the latent space. Next, we will introduce details of the new geodesic evolution layer.
@@
We utilize a multilayer neural network with $J$ hidden layers to simulate the geodesic mapping function from $z_{v_{t}}$ to $z_{v_{t+1}}$. To simplify the math notation, we let $u^{j}$ denote the input of $j$-th hidden layer and $u^{j+1}$ denotes the output, where $j\in\{1,\cdots,J\}$. A composition of a local linear transformation, $W^{j}$, and a global convolutional kernel, $\mathcal{H}^{j}$, is used to extract both global and local representations at each hidden layer. Additionally, a nonlinear activation function, $\sigma(\cdot)$, is carefully designed to encourage a smooth evolution of the geodesic path. We update $u^{j}\mapsto u^{j+1}$ by
@@
where * represents a convolution operator. In order to compute the global convolution in an efficient way, we apply matrix-vector multiplication for two Fourier signals in the frequency domain, and then project it back to the original space. We employ a Gaussian Error Linear Unit (GELU) (Hendrycks and Gimpel, 2016) with the smoothing operator $K$ to ensure the smoothness of the output signal, we have $\sigma:=K(\text{GELU}(\cdot))$.
Finally, we update $z_{v_{t-1}}$ to $z_{v_{t}}$ by
@@
We formulate the geodesic loss function of GNO as an empirical data loss term guided by numerical solutions (e.g., using an Euler integrator) to the geodesic shooting equations. Since our model GNO is now learning a sequence of latent velocity fields $\{z_{v_{0}},z_{v_{1}},\cdots,z_{v_{\tau}}\}$, the registration network will decode a time-sequence of $\{v_{0},v_{1},\cdots,v_{\tau}\}$ accordingly.
Given a set of $N$ observations, $\bm{\hat{v}}^{n}\triangleq\{\hat{v}_{1}^{n},\cdots,\hat{v}_{\tau}^{n}\}$, which are numerical solutions to the geodesic shooting equation (Eq. (2)) with a given initial condition, we formulate our geodesic loss as the mean squared error (MSE) between the predicted time-dependent velocities and the corresponding velocities computed via the numerical solution to the EPDiff equation, i.e.
$l(\theta_{r},\theta_{v})$ $=\sum_{n=1}^{N}\|\mathcal{D}_{\theta_{v}}(\mathcal{G}_{\theta_{r}}(z_{v_{0}}))-\bm{\hat{v}}^{n}\|_{2}^{2}$
$+\text{Reg}(\theta_{r},\theta_{v}),$ (5)
where $\mathcal{G}_{\theta_{r}}(z_{v_{0}})$ denotes the geodesic mapping function that transforms the latent initial velocity $z_{v_{0}}$ into a sequence of time-dependent latent velocities $z_{v_{1}},\ldots,z_{v_{\tau}}$}. The decoder $\mathcal{D}_{\theta_{v}}(\cdot)$ then maps this latent trajectory back to the image space, reconstructing the deformation process. The term $\text{Reg}(\cdot)$ represents a regularization applied to the network parameters.